# Constructing a T-test for Value Function Comparison of Individualized Treatment Regimes in the Presence of Multiple Imputation for Missing Data

### v1

@@ 1 Introduction
In these applications, decision-makers may seek to compare the effectiveness of an ITR against, for example, the observed treatment assignment, a one-size-fits-all treatment approach, or another ITR. This comparison is valuable for both patients and healthcare providers in directing their efforts toward the most effective treatment option. For example, individuals with or at risk for hypertension are generally advised to increase their physical activity. The American Heart Association, based on the Physical Activity Guidelines for Americans (Piercy et al., 2018), recommends that adults engage in 2.5-5 hours of moderate-intensity physical activity per week; although it is noted that additional health benefits can be achieved by achieving beyond 5 hours per week (Olson et al., 2019). Nevertheless, this physical activity recommendation can be challenging to meet for many people. In addition, it is possible that this level of physical activity may not be beneficial for all individuals (Bouchard and Rankinen, 2001).
@@
Thus, it is important to examine whether physical activity, especially at high recommended levels, which might be challenging to implement, uniformly benefits hypertension prevention and management across all demographic groups. This may be particularly relevant in countries in which the prevalence of anti-hypertensive medications is low, such as China. We aim to use ITRs to differentiate between subpopulations that experience incremental benefits from increasing their weekly physical activity to more (versus less or equal) than 5 hours per week and directly compare the ITR generated to population-level recommendations. For subpopulations who are unable to adhere to recommended activity levels, there are potential significant benefits conferred by focusing on alternative behavior interventions, such as dietary modification. Moreover, there are various methods for estimating the ITRs, such as Q-learning (Qian and Murphy, 2011) and D-learning (Tian et al., 2014). Direct comparisons of these ITRs are valuable for determining whether one ITR results in a superior population-level improvement, as compared to another ITR, or to a non-individualized treatment.
@@
Observational data is frequently encountered in research studies and serves as a vital data source for deriving ITRs. Observational epidemiologic studies have several potential strengths, including increased representation of variability in the target population, through large sample sizes and sampling strategies to increase population representation. However, observational data generally lack balance in treatment assignments. To address this confounding, the average outcome for the population, assuming adherence to a specific treatment regime, is weighted by inverse individual propensity scores, and this weighted outcome is defined as the value function. Researchers have commonly employed the value function to evaluate the performance of an ITR derived from observational cohort data. The presence of missing data constitutes a common challenge encountered within randomized trial datasets as well as observational datasets. While large observation data potentially provides more heterogeneity than smaller, more controlled studies, observational studies are often more susceptible to missing data problems. One efficacious approach involves the application of multiple imputation Sterne et al. (2009) whereby multiple plausible imputed datasets are generated. Each imputed dataset entails the substitution of missing values with imputed values derived from a predictive model. Subsequently, standard statistical methods are applied to each of the imputed datasets separately. This process yields a collection of results, and the observed variability among these results across the imputed datasets reflects the inherent uncertainty associated with missing values. Multiple imputation allows comprehensive assessment of the robustness of the study findings in the presence of missing data and has the potential to enhance reliability and generalizability of the research outcomes.
@@
Various approaches have been employed to estimate the variance of a value function for ITRs, such as jackknife (Jiang et al., 2021), cross-validation (Cui et al., 2017; Zhao et al., 2012), bootstrapping (Efron, 1992), and Q-function model-based approaches (Shi et al., 2022). Each is explained below. However, these approaches tend to become intricate when using multiple imputation and are not intended for directly comparing the value functions of a pair of ITRs, such as estimating the variance around the difference between two value functions. Jiang et al. (2021) employed the jackknife or leave-one-out cross-validation method to estimate the variance of the value function and compared 24 individualized treatment regimes derived from 24 machine learning models. This jackknife approach consistently estimates the variance around a value function estimate and requires only weak assumptions such as requiring the samples to be independent and identically distributed, and that as the sample size becomes larger, the decision rule estimated from n-1 individuals converges to the decision rule estimated from n individuals. However, it is a time-consuming implementation and better suited for small datasets. Cross-validation provides an alternative means to estimate variance. Cui et al. (2017) and Zhao et al. (2012) used cross-validation in their paper. The data is divided into $K$ folds, and the ITR was estimated based on the $K-1$ folds and the value function was evaluated based on the remaining fold. The process is repeated $K$ times to get $K$ estimates of value functions and then the variance of value function is calculated. A common choice of $K$ is $K=10$ which is called 10-fold cross-validation. This method is less time-consuming than the jackknife, though it has a tendency to understate the actual variance. This is because each data point is utilized in both the training and testing sets, leading to a correlation among the accuracy measures of each fold (Bates et al., 2021). Bootstrapping provides an alternative approach to compute the variance of the value function for an ITR. The standard n-out-of-n bootstrap method Efron (1992) involves randomly selecting n observations with replacements from the original data with n observations to create new datasets, which are then repeatedly generated 500 or 1000 times. Then, methods for finding the value function of the ITR are applied to each sequence, resulting in slightly different value functions. Subsequently, the variance and confidence interval of the ITR’s value function are developed. Variations of the standard n-out-of-n bootstrap method include double bootstrap Chakraborty et al. (2010), adaptive bootstrap Laber and Murphy (2011), and m-out-of-n bootstrap Chakraborty et al. (2013). The bootstrapping method requires repeating operations on the data, usually 500 times or more, making the approach computationally intensive. Shi et al. (2022) have developed a method for constructing the statistical inference of a policy’s value function for reinforcement learning when either the number of decision points or the sample trajectories diverge to infinity. Their approach involves utilizing the sieve method to approximate the Q- function and employs “SequentiAl Value Evaluation” to split data and iteratively find the optimal policy and the value function estimate. Their method constructs a valid confidence interval around the value function estimate that achieves the nominal coverage. However, it requires estimation of the Q-function and relies on the correctness of the Q-function model.
@@
In this paper, we present a new method that enables the direct comparison of any two ITRs via the value function and also provides a t-test-based p-value for the significance of observed differences. Our approach addresses the shortcomings described above in that: 1) it is less time-consuming to implement than the jackknife or bootstrapping; 2) it circumvents the problem of correlation in each fold experienced by cross-validation; and 3) it does not rely on the estimation of the Q-function. Our method is suitable for both observational studies and clinical trials. Moreover, our method enables deriving variance estimates from multiple imputed datasets without the need for additional replication or bootstrapping, making it particularly advantageous and efficient when dealing with missing data. Specifically, our approach provides a valid estimate of the variability surrounding both the value function itself and the difference between the two value functions. With our approach, it is possible to assess whether the optimal ITR significantly outperforms the one-size-fits-all approach or another ITR estimated using a different method. Furthermore, our method is characterized by its ease of understanding and implementation, while providing theoretical guarantees for estimator consistency and inference validity, including variance and confidence interval calculations. For illustrative purposes, we use two models, Q-learning (Qian and Murphy, 2011) and D-learning (Tian et al., 2014) for ITR comparison, although more sophisticated models can also be incorporated.
@@ 2 Methods
We assume that the dataset is divided into a training set that estimates the decision rule and a test set that evaluates this rule on new data, with $n$ and $m$ individuals respectively. The training data is independent of the test data. The individual index is denoted by $i$, where $X_{i}$ represents the covariate vector, $A_{i}$ the binary treatment assignment, and $Y_{i}$ the outcome. Let $\hat{d}_{1,n}$ and $\hat{d}_{2,n}$ be two decision rules we want to compare. Each $\hat{d}_{n}$ is a map from covariates to treatments. The $\hat{d}_{1,n}$ and $\hat{d}_{2,n}$ can be zero-ordered decision rules (e.g. assign everyone the same treatment), or ITRs as a function of the covariates based on models from the training set (e.g. assign everyone who is younger than 40 years old one treatment, and assign placebo otherwise). The variable $n$ indicates that the rules are estimated using training data, and in case of missing data, decision rules are derived from single or multiple imputation of the training data. Our objective is to estimate:
@@
  • The variance associated with the value function $Var(V_{m}(d_{j,n}))$
@@
  • The variance of the difference between the value functions of two decision rules $Var(V_{m}(d_{1,n})-V_{m}(d_{2,n}))$
@@ 2.1 Propensity Score
Although randomized trial data are ideal for estimating the optimal ITR, many researchers only have access to observational data. When using observational data, the confounding effect is a main concern and can be mitigated by propensity score modeling (Austin, 2011). The propensity score, denoted as $\pi(A|X)=P(A|X)$, models the probability of receiving treatment $A$ given covariate vector $X$. By weighting each individual with the inverse propensity score, the observational data can approximate the data from a randomized trial. Let $\hat{\pi}_{n}(a|x)$ be the estimated propensity score from the training set, and assume $\hat{\pi}_{n}(a|x)=\pi(a|x,\hat{\theta}_{n})$, where $\hat{\theta}_{n}$ are parameters for the propensity score model estimated based on the training set. Let $\theta_{0}$ represent the true parameter vector.
@@
$\sqrt{n}(\pi(a|x,\hat{\theta}_{n})-\pi(a|x,\theta_{0}))=\sqrt{n}(\hat{\theta}_{n}-\theta_{0})^{T}\phi_{0}(a,x)+o_{p}(1),$ (1)
[…]
@@
In general, most generalize linear models satisfy both assumptions under regularity conditions. For example, the first assumption is automatically satisfied for logistic regression. Suppose $\pi(a|x,\theta_{0})=\frac{e^{a\theta_{0}^{T}x}}{1+e^{\theta_{0}^{T}x}}$, for $a\in\{0,1\}$. The left hand side of the equation 1 becomes $\sqrt{n}(\frac{e^{a\hat{\theta}_{n}^{T}x}}{1+e^{\hat{\theta}_{n}^{T}x}}-\frac{e^{a\theta_{0}^{T}x}}{1+e^{\theta_{0}^{T}x}})$. By using the Taylor expansion for $\hat{\theta}_{n}$ at $\theta_{0}$, with $a=1$ and $a=0$, the left hand side $=\sqrt{n}(\hat{\theta}_{n}-\theta_{0})^{T}(2a-1)x\frac{e^{a\theta_{0}^{T}x}}{(1+e^{\theta_{0}^{T}x})^{2}}+o_{P}(1)$. Thus $\phi_{0}(a,x)=x(2a-1)\frac{e^{a\theta_{0}^{T}x}}{(1+e^{\theta_{0}^{T}x})^{2}}$ and $\hat{\phi}_{n}(a,x)=x(2a-1)\frac{e^{a\hat{\theta}_{n}^{T}x}}{(1+e^{\hat{\theta}_{n}^{T}x})^{2}}$. Assumption 1 is thus satisfied. And we can also show that the other part of our assumptions hold by standard empirical process arguments. Note that in randomized trial data, the propensity score will be fixed for binary treatment $\pi(a|x)=0.5$ and $\hat{\theta}_{n}=0$.
@@ 2.2 Value Function
We evaluate the goodness of a decision rule $\hat{d}_{j,n}$ by estimating the value function $V(\hat{d}_{j,n})=E(Y|A=\hat{d}_{j,n}(X),X)=E(\frac{Y1\{A=\hat{d}_{j,n}(X)\}}{P(A|X)})$ Qian and Murphy (2011). The value function can be interpreted as the inverse propensity score weighted average outcome if the population were to follow the decision rule $\hat{d}_{j,n}$. Specifically, we will calculate the value function based on the test set data, represented by the subscript m:
@@ 2.3 Variance for the Value Function
$\begin{split}\sqrt{m}(\hat{V}_{m}(\hat{d}_{j,n})-V_{0}(\hat{d}_{j,n}))&=\mathbb{G}_{m}(\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(X)\}}{\pi_{0}(A|X)})\\ &-\sqrt{m/n}\sqrt{n}(\hat{\theta}_{n}-\theta_{0})^{T}E(\phi_{0}(A,X)\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(x)\}}{\pi^{2}_{0}(A|X)}),\\ \end{split}$ (3)
where m denotes the empirical estimate based on the test set with $m$ individuals, $\mathbb{G}_{m}(U)=\sqrt{m}[\mathbb{P}_{m}(U)-P(U)]\rightarrow N(0,Var(U))$ by standard empirical process arguments, $\mathbb{P}_{m}(U)=m^{-1}\sum_{i=1}^{m}U_{i}$ is the empirical measure, $P(U)=E(U)$ is the expectation taken over $U$, and $\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\rightarrow N(0,\Sigma_{0})$ by Assumption 1.
Let $U_{0}=\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(x)\}}{\pi_{0}(A|X)}$ and $W_{0}=E(\phi_{0}(A,X)\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(x)\}}{\pi^{2}_{0}(A|X)})$. Standard empirical process arguments yield $\mathbb{G}_{m}(U_{0})=\sqrt{m}[\mathbb{P}_{n}(U_{0})-P(U_{0})]\rightarrow N(0,Var(U_{0}))$. We can also show that the random variables $\sqrt{m}(\hat{V}_{m}(\hat{d}_{j,n})-V_{0}(\hat{d}_{j,n}))$ converge in distribution to a mean zero normal distribution:
$\sqrt{m}(\hat{V}_{m}(\hat{d}_{j,n})-V_{0}(\hat{d}_{j,n}))\rightarrow N(0,Var(U_{0})+\frac{m}{n}W_{0}^{T}\Sigma_{0}W_{0}).$
In this context, as the value of $n$ approaches infinity, we assume the quotient $m/n$ asymptotically approaches a finite limit, rather than diverging towards infinity. Thus the expected value of the variance of a single ITR $\hat{d}_{j,n}$ is:
$Var(\hat{V}_{m}(\hat{d}_{j,n}))=m^{-1}(Var(U_{0})+\frac{m}{n}W_{0}^{T}\Sigma_{0}W_{0})=m^{-2}\sum_{i=1}^{m}(\hat{U}_{i,j}-\bar{U}_{i,j})^{2}+\frac{1}{n}W_{0}^{T}n\hat{\Sigma}_{n}W_{0}.$
The variance $\sigma_{0}^{2}$ for the value function $V_{m}(d_{j})$ is then estimated by:
$\hat{\sigma}^{2}_{m,j}=Var(\hat{V}_{m}(\hat{d}_{j}))=m^{-2}\sum_{i=1}^{m}(\hat{U}_{i,j}-\bar{U}_{i,j})^{2}+\frac{1}{n}\hat{W}_{j,m}^{T}n\hat{\Sigma}_{n}\hat{W}_{j,m}.$ (4)
Note that the two parts in the variance equation are dependent. When the number of individuals in the test set is much less than that in the training set, then we can ignore the second term $\frac{1}{n}\hat{W}_{j,m}^{T}n\hat{\Sigma}_{n}\hat{W}_{j,m}$. When we use randomized trial data, we don’t need to estimate the propensity score, and thus $\Sigma_{n}=Var(\hat{\theta}_{n})=0$, and hence only the first term remains. The estimated influence function for the training set and estimate for $W_{0}$ are as follows:
$\hat{U}_{i,j}=(Y_{i}-\hat{V}_{n}(\hat{d}_{j,n}))\frac{1\{A_{i}=\hat{d}_{j,n}\}}{\hat{\pi}_{n}(A_{i}|X_{i})},\quad i=1,...,n,$ (5)
$\hat{W}_{j,n,}=n^{-1}\sum_{i=1}^{n}[\frac{\hat{\phi}(A_{i},X_{i})\hat{U_{i,j}}}{\hat{\pi}_{n}(A_{i}|X_{i})}],$ (6)
where $\hat{\phi}_{n}(A_{i},X_{i})=x_{i}(2a_{i}-1)\frac{e^{a_{i}\hat{\theta}_{n}^{T}x_{i}}}{(1+e^{\hat{\theta}_{n}^{T}x_{i}})^{2}}$, $\bar{U}_{j,m}=m^{-1}\sum_{i=1}^{m}\hat{U}_{i,j}$ (j=1,2), and $\hat{\Sigma}_{n}=Var(\hat{\theta}_{n})$ is estimated from the training set based on the variance due to modeling the propensity score.

### v3

@@ 1 Introduction
In these applications, decision-makers may seek to compare the effectiveness of an ITR against, for example, the observed treatment assignment, a one-size-fits-all treatment approach, or another ITR. This comparison is valuable for both patients and healthcare providers in directing their efforts toward the most effective treatment option. For example, individuals with or at risk for hypertension are generally advised to increase their physical activity. The American Heart Association, based on the Physical Activity Guidelines for Americans 1, recommends that adults engage in 2.5-5 hours of moderate-intensity physical activity per week; although it is noted that additional health benefits can be achieved by achieving beyond 5 hours per week 2. Nevertheless, this physical activity recommendation can be challenging to meet for many people. In addition, it is possible that this level of physical activity may not be beneficial for all individuals 3.
@@
Thus, it is important to examine whether physical activity, especially at high recommended levels, which might be challenging to implement, uniformly benefits hypertension prevention and management across all demographic groups. This may be particularly relevant in countries in which the prevalence of anti-hypertensive medications is low, such as China. We aim to use ITRs to differentiate between subpopulations that experience incremental benefits from increasing their weekly physical activity to more (versus less or equal) than 5 hours per week and directly compare the ITR generated to population-level recommendations. For subpopulations who are unable to adhere to recommended activity levels, there are potential significant benefits conferred by focusing on alternative behavior interventions, such as dietary modification. Moreover, there are various methods for estimating the ITRs, such as Q-learning and D-learning 4, 5. Direct comparisons of these ITRs are valuable for determining whether one ITR results in a superior population-level improvement, as compared to another ITR, or to a non-individualized treatment.
@@
Observational data is frequently encountered in research studies and serves as a vital data source for deriving ITRs. Observational epidemiologic studies have several potential strengths, including increased representation of variability in the target population, through large sample sizes and sampling strategies to increase population representation. However, observational data generally lack balance in treatment assignments. To address this confounding, the average outcome for the population, assuming adherence to a specific treatment regime, is weighted by inverse individual propensity scores, and this weighted outcome is defined as the value function. Researchers have commonly employed the value function to evaluate the performance of an ITR derived from observational cohort data. The presence of missing data constitutes a common challenge encountered within randomized trial datasets as well as observational datasets. While large observation data potentially provides more heterogeneity than smaller, more controlled studies, observational studies are often more susceptible to missing data problems. One efficacious approach involves the application of multiple imputation 6 whereby multiple plausible imputed datasets are generated. Each imputed dataset entails the substitution of missing values with imputed values derived from a predictive model. Subsequently, standard statistical methods are applied to each of the imputed datasets separately. This process yields a collection of results, and the observed variability among these results across the imputed datasets reflects the inherent uncertainty associated with missing values. Multiple imputation allows comprehensive assessment of the robustness of the study findings in the presence of missing data and has the potential to enhance reliability and generalizability of the research outcomes.
@@
Various approaches have been employed to estimate the variance of a value function for ITRs, such as jackknife7, cross-validation 8, 9, Q-function model-based approaches 10, and bootstrapping 11. Each is explained below. However, these approaches tend to become intricate when using multiple imputation and are not intended for directly comparing the value functions of a pair of ITRs, such as estimating the variance around the difference between two value functions. Jiang et al7. employed the jackknife or leave-one-out cross-validation method to estimate the variance of the value function and compared 24 individualized treatment regimes derived from 24 machine learning models. This jackknife approach consistently estimates the variance around a value function estimate and requires only weak assumptions such as requiring the samples to be independent and identically distributed, and that as the sample size becomes larger, the decision rule estimated from n-1 individuals converges to the decision rule estimated from n individuals. However, it is a time-consuming implementation and better suited for small datasets. Cross-validation provides an alternative means to estimate variance. Cui et al.8 and Zhao et al. 9 used cross-validation in their paper. The data is divided into $K$ folds, and the ITR was estimated based on the $K-1$ folds and the value function was evaluated based on the remaining fold. The process is repeated $K$ times to get $K$ estimates of value functions and then the variance of value function is calculated. A common choice of $K$ is $K=10$ which is called 10-fold cross-validation. This method is less time-consuming than the jackknife, though it has a tendency to understate the actual variance. This is because each data point is utilized in both the training and testing sets, leading to a correlation among the accuracy measures of each fold. 12 Shi et al.10 have developed a method for constructing the statistical inference of a policy’s value function for reinforcement learning when either the number of decision points or the sample trajectories diverge to infinity. Their approach involves utilizing the sieve method to approximate the Q- function and employs “SequentiAl Value Evaluation” to split data and iteratively find the optimal policy and the value function estimate. Their method constructs a valid confidence interval around the value function estimate that achieves the nominal coverage. However, it requires estimation of the Q-function and relies on the correctness of the Q-function model. Bootstrapping provides an alternative approach to compute the variance of the value function for an ITR. The standard n-out-of-n bootstrap method 11 involves randomly selecting n observations with replacements from the original data with n observations to create new datasets, which are then repeatedly generated 500 or 1000 times. Then, methods for finding the value function of the ITR are applied to each sequence, resulting in slightly different value functions. Subsequently, the variance and confidence interval of the ITR’s value function are developed. Variations of the standard n-out-of-n bootstrap method include double bootstrap13, adaptive bootstrap14, and m-out-of-n bootstrap 15. All bootstrapping methods require repeating operations on the data, usually 500 times or more, making the approach computationally intensive.
@@
In this paper, we present a new method that enables the direct comparison of any two ITRs via the value function and also provides a t-test-based p-value for the significance of observed differences. Our approach addresses the shortcomings described above in that: 1) it is less time-consuming to implement than the jackknife or bootstrapping; 2) it circumvents the problem of correlation in each fold experienced by cross-validation; and 3) it does not rely on the estimation of the Q-function. Our method is suitable for both observational studies and clinical trials. Moreover, our method enables deriving variance estimates from multiple imputed datasets without the need for additional replication or bootstrapping, making it particularly advantageous and efficient when dealing with missing data. Specifically, our approach provides a valid estimate of the variability surrounding both the value function itself and the difference between the two value functions. With our approach, it is possible to assess whether the optimal ITR significantly outperforms the one-size-fits-all approach or another ITR estimated using a different method. Furthermore, our method is characterized by its ease of understanding and implementation, while providing theoretical guarantees for estimator consistency and inference validity, including variance and confidence interval calculations. For illustrative purposes, we use two models, Q-learning 4 and D-learning 5 for ITR comparison, although more sophisticated models can also be incorporated.
@@ 2 Methods
In this section, we detail the methodology for calculating the t-test based p-value to assess the difference between two individualized treatment rule (ITR) value functions. We initiate our discussion by defining the framework and introducing relevant notations. Subsequently, we present Proposition 1, which outlines the limiting distribution of the variance estimation for the estimated value function. We then proceed to establish the limiting distribution of the difference between two value functions as delineated in Proposition 2, and describe the derivation of the corresponding t-test statistics for pairwise value function comparisons in Equation 12. Lastly, we discuss the adaptation of this variance estimation and t-test statistics to accommodate datasets with missing entries via multiple imputation in Equation 15.
We assume that the dataset is divided into a training set that estimates the decision rule and a test set that evaluates this rule on new data, with $n$ and $m$ individuals respectively. The training data is independent of the test data. The individual index is denoted by $i$, where $X_{i}$ represents the covariate vector, $A_{i}$ the binary treatment assignment, and $Y_{i}$ the outcome. Let $\hat{d}_{1,n}$ and $\hat{d}_{2,n}$ be two decision rules we want to compare. Each $\hat{d}_{n}$ is a map from covariates to treatments. The $\hat{d}_{1,n}$ and $\hat{d}_{2,n}$ can be zero-ordered decision rules (e.g. assign everyone the same treatment), or ITRs as a function of the covariates based on models from the training set (e.g. assign everyone who is younger than 40 years old one treatment, and assign placebo otherwise). The variable $n$ indicates that the rules are estimated using training data, and in case of missing data, decision rules are derived from single or multiple imputation of the training data. The main contribution of our new method provide the estimates for the following:
@@
  • The variance (and 95% confidence interval) associated with the value function $Var(V_{m}(d_{j,n}))$
@@
  • The variance (and 95% confidence interval) of the difference between the value functions of two decision rules $Var(V_{m}(d_{1,n})-V_{m}(d_{2,n}))$
@@ 2.1 Propensity Score
Although randomized trial data are ideal for estimating the optimal ITR, many researchers only have access to observational data. When using observational data, the confounding effect is a main concern and can be mitigated by propensity score modeling 16. The propensity score, denoted as $\pi(A|X)=P(A|X)$, models the probability of receiving treatment $A$ given covariate vector $X$. By weighting each individual with the inverse propensity score, the observational data can approximate the data from a randomized trial. Let $\hat{\pi}_{n}(a|x)=\pi_{n}(a|x,\hat{\theta}_{n})$ be the estimated propensity score from the training set, and $\pi_{0}(a|x)=\pi_{n}(a|x,\theta_{0})$ be the true propensity score, where $\hat{\theta}_{n}$ and $\theta_{0}$ denote the estimated and the true parameters for the propensity score model.
@@
Let $\phi_{0}(A,X)$ be a vector of functions of $A$ and $X$, such that the following equation is satisfied:
$\sqrt{n}(\pi(a|x,\hat{\theta}_{n})-\pi(a|x,\theta_{0}))=\sqrt{n}(\hat{\theta}_{n}-\theta_{0})^{T}\phi_{0}(a,x)+o_{p}(1).$ (1)
Assumption 2.
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\rightarrow N(0,\Sigma_{0}),$ (2)
and there exists a consistent estimate $\hat{\Sigma}_{n}$ for $\Sigma_{0}$.
@@
In general, most generalize linear models satisfy all assumptions under regularity conditions. The assumptions can be verified using a Taylor expansion, the asymptotic normality of the maximum likelihood estimator, and standard empirical process arguments. For illustration, detailed derivations for the logistic regression case are provided in Appendix A. Note that in randomized trial data, the propensity score will be fixed for binary treatment $\pi(a|x)=0.5$ and $\hat{\theta}_{n}=0$.
@@ 2.2 Value Function
We evaluate the goodness of a decision rule $\hat{d}_{j,n}$ by estimating the value function $V(\hat{d}_{j,n})=E(Y|A=\hat{d}_{j,n}(X),X)=E(\frac{Y1\{A=\hat{d}_{j,n}(X)\}}{P(A|X)})$ 4. The value function can be interpreted as the inverse propensity score weighted average outcome if the population were to follow the decision rule $\hat{d}_{j,n}$. Specifically, we will calculate the value function based on the test set data, represented by the subscript m:
@@ 2.3 Variance for the Value Function
Let $Z=(X,A,Y)$. For any measurable function of $Z$, let $U=U(Z)$, $\mathbb{P}_{m}(U)=m^{-1}\sum_{i=1}^{m}U_{i}$ be the empirical measure, and $P(U)=E(U)$ be the expectation taken over $U$. Then $\mathbb{G}_{m}(U)=\sqrt{m}[\mathbb{P}_{m}(U)-P(U)]\rightarrow N(0,Var(U))$ by standard empirical process arguments. Through some derivations we can derive the following equation (detailed proofs are attached in Appendix A):
$\begin{split}\sqrt{m}(\hat{V}_{m}(\hat{d}_{j,n})-V_{0}(\hat{d}_{j,n}))&=\mathbb{G}_{m}(\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(X)\}}{\pi_{0}(A|X)})\\ &-\sqrt{m/n}\sqrt{n}(\hat{\theta}_{n}-\theta_{0})^{T}E(\phi_{0}(A,X)\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(X)\}}{\pi^{2}_{0}(A|X)})+o_{P}(1),\\ \end{split}$ (4)
where m denotes the empirical estimate based on the test set with $m$ individuals.
Let $U=U(Z)=\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(X)\}}{\pi_{0}(A|X)}$ and $W=W(Z)=E(\phi_{0}(A,X)\frac{(Y-V_{0}(\hat{d}_{j,n}))1\{A=\hat{d}_{j,n}(X)\}}{\pi^{2}_{0}(A|X)})$. We can show that the following proposition holds.
@@
$\sqrt{m}(\hat{V}_{m}(\hat{d}_{j,n})-V_{0}(\hat{d}_{j,n}))\rightarrow N(0,Var(U)+\frac{m}{n}W^{T}\Sigma_{0}W),$ (5)
provided that the quotient $m/n$ asymptotically approaches a finite limit.
Standard empirical process arguments yield $\mathbb{G}_{m}(U)=\sqrt{m}[\mathbb{P}_{n}(U)-P(U)]\rightarrow N(0,Var(U))$. Since $\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\rightarrow N(0,\Sigma_{0})$, we can also show that the random variables $\sqrt{m}(\hat{V}_{m}(\hat{d}_{j,n})-V_{0}(\hat{d}_{j,n}))$ converge in distribution to a mean zero normal distribution. In this context, as the value of $n$ approaches infinity, we assume the quotient $m/n$ asymptotically approaches a finite limit, rather than diverging towards infinity.
where the estimated influence function for $U$ and $W$ are as follows:
$\hat{U}_{i,j}=(Y_{i}-\hat{V}_{n}(\hat{d}_{j,n}))\frac{1\{A_{i}=\hat{d}_{j,n}\}}{\hat{\pi}_{n}(A_{i}|X_{i})},$ (6)
$\bar{U}_{j}=m^{-1}\sum_{i=1}^{m}\hat{U}_{i,j},$ (7)
$\hat{W}_{j}=m^{-1}\sum_{i=1}^{m}[\frac{\hat{\phi}(A_{i},X_{i})\hat{U}_{i,j}}{\hat{\pi}_{n}(A_{i}|X_{i})}],$ (8)
where $i=1,...,n$.
Thus, the variance $\sigma_{0}^{2}$ for the value function $V_{m}(d_{j})$ can be estimated consistently by:
$\hat{\sigma}^{2}_{m,j}=Var(\hat{V}_{m}(\hat{d}_{j}))\approx m^{-2}\sum_{i=1}^{m}(\hat{U}_{i,j}-\bar{U}_{j})^{2}+\frac{1}{n}\hat{W}_{j}^{T}\hat{\Sigma}_{n}\hat{W}_{j}.$ (9)
Note that the two parts in the variance equation are dependent. When the number of individuals in the test set is much less than that in the training set ($m<<n$), then we can ignore the second term $\frac{1}{n}\hat{W}_{j}^{T}\hat{\Sigma}_{n}\hat{W}_{j}$. [^1] When the number of individuals in the test set is much less than that in the training set, $m<<n$. Thus $m^{-1}>>n^{-1}$. Looking at equation (9), $m^{-2}\sum_{i=1}^{m}(\hat{U}_{i,j}-\bar{U}_{j})^{2}>>\frac{1}{n}\hat{W}_{j}^{T}\hat{\Sigma}_{n}\hat{W}_{j}$. When we use randomized trial data, we don’t need to estimate the propensity score, and thus $\hat{\Sigma}_{n}=0$ in this situation, and hence only the first term remains.